Circular No. 5767

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

NOVA OPHIUCHI 1993 M. Della Valle, European Southern Observatory; and M. Turatto and

M. Lazzarin, Osservatorio Astronomico di Padova, report: “Spectrograms (range 400-1000 nm, resolution ∼ 1 nm) obtained on Apr. 18.4 UT, with the New Technology Telescope (+ EMMI) and the 1.52-m reflector (+ Boller & Chivens spectrograph) at La Silla, confirm this object as a nova caught during its early decline. The spectrum is dominated by emission lines of H, Fe II (mult. 42, 49, 73, 74), Na I, N I, O I, and Ca II. Hα is flanked by a weak, blue-shifted, P-Cyg profile. The expansion velocity derived from the minimum of the P-Cyg absorption is ∼ 3600 km/s.”

Further visual magnitude estimates: Apr. 17.51 UT, 9.7 (A. Jones, Nelson, New Zealand); 17.52, 9.4 (D. A. J. Seargent, The Entrance, New South Wales, Australia); 17.74, 9.4 (Jones); 18.54, 9.5 (P. F. Williams, Heathcote, N.S.W.); 18.70, 9.4 (Y. Kushida, Yatsugatake South Base Observatory; photoelectric V); 18.70, 9.3 (S. Takahashi, Shiga, Japan); 18.81, 9.4 (P. J. Camilleri, Cobram, Vic., Australia).

SUPERNOVA 1993J IN NGC 3031 L. V. Morrison and R. W. Argyle, Royal Greenwich Observatory; and

L. Helmer, Copenhagen University Observatory, report the following accurate optical position, obtained with the Carlsberg Automatic Meridian Circle on La Palma: α = 9^{h}55^{m}24^{s}.778, δ = +69^{o}01′13′′.70 (equinox J2000.0, FK5 reference frame; epoch 1993.27; accuracy 0′′.10 in α and δ); the corresponding B1950.0 position is α = 9^{h}51^{m}19^{s}.045, δ = +69^{o}15′26′′.42. The position is the mean of seven observations, and will be improved further after more observations are secured.

Further photoelectric photometry (cf. IAUC 5750, 5755, 5758): Apr. 2.93 UT, V = 11.42 (Morrison, Argyle, and Helmer; same instrument as used above; ± 0.04); 7.92, V = 11.50 (Morrison et al.); 8.91, V = 11.47 (Morrison et al.); 10.91, V = 11.23 (Morrison et al.); 15.90, V = 10.91 (Morrison et al.); 16.88, V = 10.80 ± 0.04, B–V = +0.59 ± 0.05 (D. Hanzl, N. Copernicus Planetarium, Brno); 16.90, V = 10.88 (Morrison et al.); 17.86, V = 10.81 ± 0.03, B–V = +0.63 ± 0.03 (Hanzl); 17.90, V = 10.81 (Morrison et al.).

1993 April 19 Daniel W. E. Green